### v1

# LLM-Meta-SR: Learning to Evolve Selection Operators for Symbolic Regression

Abstract
Large language models (LLMs) have revolutionized algorithm development, yet their application in symbolic regression, where algorithms automatically discover symbolic expressions from data, remains constrained and is typically designed manually by human experts. In this paper, we propose a learning-to-evolve framework that enables LLMs to automatically design selection operators for evolutionary symbolic regression algorithms. We first identify two key limitations in existing LLM-based algorithm evolution techniques: code bloat and a lack of semantic guidance. Bloat results in unnecessarily complex components, and the absence of semantic awareness can lead to ineffective exchange of useful code components, both of which can reduce the interpretability of the designed algorithm or hinder evolutionary learning progress. To address these issues, we enhance the LLM-based evolution framework for meta symbolic regression with two key innovations: bloat control and a complementary, semantics-aware selection operator. Additionally, we embed domain knowledge into the prompt, enabling the LLM to generate more effective and contextually relevant selection operators. Our experimental results on symbolic regression benchmarks show that LLMs can devise selection operators that outperform nine expert-designed baselines, achieving state-of-the-art performance. This demonstrates that LLMs can exceed expert-level algorithm design for symbolic regression.
1 Introduction
Symbolic regression is the task of discovering mathematical expressions that accurately model a given dataset. It is widely used in domains such as finance [54], education [53],chemistry [9], and physics [15], due to its interpretability and ability to uncover underlying relationships. Formally, given a dataset $\mathcal{D}=\{\mathbf{x}^{(i)},y^{(i)}\}_{i=1}^{N}$, where $\mathbf{x}^{(i)}\in\mathbb{R}^{d}$ represents input features and $y^{(i)}\in\mathbb{R}$ denotes the corresponding target value, the goal is to find a symbolic expression $f^{*}\in\mathcal{F}$ from a space of candidate expressions $\mathcal{F}$ that minimizes a predefined loss function $\mathcal{L}$, i.e.,
$f^{*}=\arg\min_{f\in\mathcal{F}}\mathcal{L}(f(\mathbf{x}),y).$ (1)
Numerous efforts have been devoted to designing symbolic regression algorithms. These approaches can be broadly categorized as evolutionary symbolic regression [5, 17], neural symbolic regression [3, 27], and search-based symbolic regression [65, 70] approaches. In recent years, there has been growing interest in hybrid methods that combine the flexibility of evolutionary algorithms with the learning capabilities of neural networks [32, 22, 71].
In both evolutionary [5, 17] and neural-evolutionary [32, 22, 71] symbolic regression algorithms, the system follows an iterative optimization process in which solutions are progressively refined by selecting promising candidates. In this process, the selection operator responsible for identifying candidate solutions worth refining, plays a key role in system effectiveness. However, in current evolutionary and neural-evolutionary symbolic regression systems, the selection operator is typically manually designed by domain experts. This manual process requires significant effort and trial-and-error, which can limit the pace of progress. Ideally, an automated algorithm design framework capable of generating effective selection operators for symbolic regression would substantially enhance research and development efficiency in this area.
Large language models (LLMs) have recently emerged as powerful tools for automated heuristic design in various optimization tasks [52, 35, 69]. However, most existing frameworks primarily focus on generating heuristic rules for specific combinatorial optimization tasks. In this paper, we pursue a more ambitious goal: to automatically design a core algorithmic component—specifically, the selection operator—in evolutionary symbolic regression, with the aim of achieving strong performance across a wide range of regression tasks. While frameworks such as FunSearch [52], EoH [35], and ReEvo [69] provide useful foundations for algorithm evolution, two key challenges remain underexplored in LLM-driven algorithm evolution. The first is complexity bias, illustrated in Figure 2. LLMs tend to generate overly long or intricate code during the optimization process, reflecting a phenomenon analogous to code bloat in genetic programming [2]. This results in the accumulation of redundant or non-functional logic, which impairs interpretability, wastes a large number of tokens, and slows optimization. As shown in Figure 2, such complexity can lead to performance stagnation, ultimately limiting the effectiveness of evolution.
Figure 1: Average code length of solutions in the population over generations, along with the score of the best solution.
Figure 2: Illustration of the importance of fine-grained semantic differences between solutions during algorithm evolution.
The second is the underutilization of semantic information in the generated code. Here, we define semantic information as the performance of an algorithm on individual task instances. Typically, only the average performance across all training instances is provided to the LLM to guide code generation [23]. However, this average does not sufficiently capture the fine-grained behavior of the algorithm across different tasks. As shown in Figure 2, consider a scenario with three algorithms: the first two perform well on the first dataset but poorly on the second, while the third shows the opposite pattern. Despite their contrasting behaviors, all three algorithms may exhibit similar average performance. However, presenting the first two algorithms together to the LLM offers limited benefit, as they exhibit redundant patterns—both succeeding and failing on the same datasets. In contrast, combining the first and third algorithms allows the LLM to integrate complementary strengths, potentially generating an algorithm that performs well on both datasets. Thus, incorporating semantic information is crucial for effective algorithm evolution.
In this paper, we propose an LLM-driven method for meta symbolic regression to automatically design selection operators  [^1] Source Code: https://anonymous.4open.science/r/LLM-Meta-SR/. The main contributions are as follows:
  • We propose an LLM-driven meta symbolic regression framework to automatically design selection operators. This learning-to-evolve process enables the discovery of generalizable operators that consistently outperform human-designed counterparts on a wide range of unseen datasets.
  • We identify the issue of bloat in LLM-based code generation and introduce a bloat control strategy that improves both the interpretability of the evolved code and the effectiveness of the evolutionary process.
  • We propose semantic-based feedback and complementary selection mechanisms to fully leverage semantic information during LLM-driven generation, explicitly guiding algorithm evolution by integrating effective building blocks and enhancing learning performance.
  • We design ideal properties of selection operators based on domain knowledge to craft more effective prompts that guide LLMs in generating high-quality selection operators.
2 Background and Related Work
2.1 Selection Operators in Symbolic Regression
There has been significant research on designing selection operators for symbolic regression, including tournament selection [63], Boltzmann sampling [56, 52], lexicase selection [30], novelty search [33], and complementary phenotype selection [14]. These operators aim to improve symbolic regression algorithms from different perspectives. Lexicase selection highlights the importance of specialists on subsets of training instances [30]. Novelty search emphasizes diversity by explicitly encouraging solutions that behave differently from one another to escape local optima [33]. Complementary phenotype selection focuses on leveraging complementary individuals in crossover and using this information during the selection process [14]. Despite these advances, there remains a lack of selection operators that simultaneously capture multiple desirable traits. A key reason is that designing such operators is labor-intensive and relies heavily on trial and error.
2.2 LLMs for Symbolic Regression
Early work on LLMs for symbolic regression focused on specialized language models [27, 3], often assisted by Monte Carlo Tree Search [28, 55] or evolutionary algorithms [71]. With the advent of general-purpose LLMs, their use in symbolic regression has attracted increasing attention [56], either through the FunSearch framework [52] or via integration with evolutionary algorithm frameworks [22]. In these studies, LLMs effectively replace traditional crossover and mutation operators to generate candidate solutions, but selection operators are still manually designed. This motivates the exploration of automated selection operator design using LLMs.
Some recent work incorporates dataset descriptions into symbolic regression, referring to this as “semantics” [37]. However, this differs from the concept of semantics in genetic programming [43], where it refers to the behavior of a program. In LLM-based program evolution, the fine-grained behavior of candidate programs is often overlooked and reduced to aggregate scores, which can obscure meaningful information and hinder evolutionary progress.
3 Algorithm Framework
This work introduces a learning-to-evolve framework that leverages large language models (LLMs) within a meta symbolic regression setting to automate the design of selection operators. The key components are as follows.
3.1 Solution Representation
Each solution is a piece of code representing a selection operator. The input to the selection operator is a list of individuals. Each individual contains a list representing squared error, a list of predicted values, and a list of nodes, which can be used to compute the height and depth of the symbolic tree. The evolutionary status, specifically the ratio between the current generation number and the total number of generations, is also provided. The expected output of the selection operator is a list of promising symbolic trees. The structure of the selection operator is presented in LABEL:lst:template of the supplementary material.
3.2 Meta-Evolution Workflow
Figure 3: Workflow of LLM-driven selection operator evolution.
The meta-evolution algorithm consists of two nested loops: an outer meta-evolution loop and an inner symbolic regression loop. The outer loop is responsible for generating new selection operators, while the inner loop evaluates their performance. An overview of the workflow is shown in Figure 3.
Let $\mathcal{P}^{(t)}=\{O_{1}^{(t)},O_{2}^{(t)},\dots,O_{N}^{(t)}\}$ denote the population of selection operators at generation $t$, where $N$ is the population size and each $O_{i}^{(t)}$ is a selection operator. The fitness of each operator $O_{i}^{(t)}$, denoted $f(O_{i}^{(t)})$, is evaluated based on its performance in the symbolic regression loop. The meta-evolution process includes the following components:
  • Population Initialization: The initial population $\mathcal{P}^{(0)}$ is generated by prompting the LLM to produce $N$ random selection operators. The details of the initialization prompts are provided in Appendix A of the supplementary material.
  • Synthetic Evaluation: Since LLM-generated code may contain syntax errors or infinite loops, we run code on several basic test cases to filter out invalid operators. Each $O_{i}^{(t)}\in\mathcal{P}^{(t)}$ is evaluated in a synthetic testing environment and discarded if it contains syntax/runtime errors, exceeds a 300-second execution time, or is more than 100$\times$ slower than the fastest operator in the batch. The test cases are provided in Appendix C of the supplementary material.
  • Solution Evaluation: For each selection operator $O_{i}^{(t)}\in\mathcal{P}^{(t)}$, its fitness is computed as $f(O_{i}^{(t)})=\frac{1}{T}\sum_{j=1}^{T}\texttt{SR}(O_{i}^{(t)},\mathcal{D}_{j}),$ where $\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ are symbolic regression datasets, and $\texttt{SR}(O_{i}^{(t)},\mathcal{D}_{j})$ denotes the symbolic regression performance of $O_{i}^{(t)}$ on dataset $\mathcal{D}_{j}$. In brief, each dataset is split into a training and validation set. Symbolic regression is performed on the training set, and the fitness of the selection operator is computed based on its performance on the validation set. Details of the solution evaluation procedure are provided in Appendix D of the supplementary material.
  • Survival Selection: A subset of operators is retained to form the next generation $\mathcal{P}^{(t+1)}$. To control bloat, a multi-objective survival selection strategy is employed, as described in Section 3.3.
  • Solution Selection: Two parent operators $O_{a}^{(t)},O_{b}^{(t)}\in\mathcal{P}^{(t)}$ are selected for recombination. The selection strategy incorporates semantic information, as detailed in Section 3.4.
  • Solution Generation: Based on the selected parents, $N$ candidate operators are generated per generation. Of these, $N-M$ are generated via crossover and $M$ via mutation:
    • Operator Crossover: For each pair $(O_{a}^{(t)},O_{b}^{(t)})$, a new operator is generated as $O_{\text{new}}^{(t+1)}=\texttt{LLM\_Crossover}(O_{a}^{(t)},O_{b}^{(t)})$. The goal is to generate a new piece of code that combines effective building blocks from the selected parent operators, guided by their performance scores on the evaluated datasets.
    • Operator Mutation: Let $O^{*}=\arg\max_{O_{i}^{(t)}\in\mathcal{P}^{(t)}}f(O_{i}^{(t)})$ be the best-performing operator in generation $t$ [69]. Mutated variants are generated as $O_{\text{mut}}^{(t+1)}=\texttt{LLM\_Mutate}(O^{*})$, with the objective of generating novel code based on the elite operator.
    The prompts used for crossover and mutation are provided in Appendix A.
This process continues until a predefined number of generations $T_{\text{max}}$ is reached. The operator with the highest fitness value across all generations is selected as the final selection operator.
3.3 Bloat Control
Prompt-based Length Limit: To mitigate code bloat in the evolution of selection operators, we incorporate a length constraint in the prompt during solution generation. For any operator $O_{i}\in\mathcal{P}^{(t)}$, its code length is denoted by $\ell(O_{i})$, measured as the number of non-empty, non-comment lines in the implementation. Rather than enforcing a hard upper bound $\ell(O_{i})\leq L_{\max}$, which is difficult to control during LLM generation, we embed the maximum length $\ell_{\text{target}}$ directly into the prompt. For instance, the LLM is instructed: “Write a selection operator with code length $\leq\ell_{\text{target}}$.” This prompt-guided strategy encourages the model to produce more concise programs.
Multi-Objective Survival Selection: Beyond prompt-based constraints, we employ a multi-objective survival selection strategy based on both operator fitness and code length to further control bloat, since LLMs do not always follow instructions to generate code within a specified length. Each operator $O_{i}$ is represented as a tuple $(f(O_{i}),\ell(O_{i}))$, where $f(O_{i})$ denotes its average task performance across $T$ tasks, and $\ell(O_{i})$ denotes its code length. The parents for generating the next generation $\mathcal{P}^{(t+1)}$ are selected from the combined set $\mathcal{P}^{(t)}\cup\mathcal{P}^{\text{offspring}}$ using a dominance-dissimilarity selection mechanism [67].
The key idea is to compute a dominance score for each operator based on weak Pareto dominance [34] and code similarity. An operator $O_{i}$ is said to weakly Pareto dominate another operator $O_{j}$, denoted $O_{i}\succeq O_{j}$, if and only if $f(O_{i})\geq f(O_{j})$ and $\ell(O_{i})\leq\ell(O_{j})$. For each such pair $(O_{i},O_{j})$, the score of $O_{j}$ is penalized by the code similarity $\text{sim}(O_{i},O_{j})$, computed using the CodeBLEU metric [51]. The total dominance score of $O_{j}$ is then defined as $s(O_{j})=\sum_{O_{i}\succeq O_{j}}-\text{sim}(O_{i},O_{j})$. Operators are then ranked, and the top-$N$ operators with higher scores are selected, as they are less frequently dominated and exhibit greater dissimilarity from dominating counterparts. This selection strategy encourages a trade-off between maximizing task performance and minimizing code complexity, effectively reducing operator bloat while maintaining diversity and effectiveness.
3.4 Semantics-Aware Evolution
Semantic-based Selection: In the meta-evolution scenario, the goal of the crossover operator is to combine the strengths of two LLM-generated selection operators. As introduced in Section 1, crossover benefits more from combining solutions with complementary strengths than from pairing generalists.
To this end, we propose a semantics-aware selection strategy, with the pseudocode provided in Algorithm 1. Because each selection operator $O_{i}\in\mathcal{P}^{(t)}$ has been evaluated across $d$ datasets and is associated with a score vector $\mathbf{s}_{i}=[s_{i,1},\dots,s_{i,d}]$, we can compute a complementarity score for each candidate in the population. The process begins by randomly selecting the first parent $O_{a}\in\mathcal{P}^{(t)}$. Random selection, as used in ReEvo [69] and HSEvo [12], helps mitigate premature convergence [12]. Then, for each candidate $O_{i}\in\mathcal{P}^{(t)}$, we compute a complementarity score:
$\mu_{i}=\frac{1}{d}\sum_{j=1}^{d}\max(s_{a,j},s_{i,j}),$ (2)
which measures the potential combined performance of $O_{a}$ and $O_{i}$ across all datasets. The second parent $O_{b}$ is selected as the operator with the highest complementarity score, i.e., $O_{b}=O_{i^{*}}$ where $i^{*}=\arg\max_{i}\mu_{i}$.
Algorithm 1 Semantics-Aware Selection
0:  $\mathcal{P}^{(t)}=\{O_{1},\dots,O_{N}\}$: population where each $O_{i}$ has score vector $\mathbf{s}_{i}$
1:  $O_{a}\leftarrow$ Random sample from $\mathcal{P}^{(t)}$
2:  for each $O_{i}\in\mathcal{P}^{(t)}$ do
3:     Compute $\mu_{i}\leftarrow\frac{1}{d}\sum_{j=1}^{d}\max(s_{a,j},s_{i,j})$
4:  end for
5:  $i^{*}\leftarrow\arg\max_{i}\mu_{i}$
6:  $O_{b}\leftarrow O_{i^{*}}$
7:  return  $O_{b}$
Semantic Feedback: The semantic selection strategy ensures that crossover is guided by semantically diverse behaviors. Furthermore, to enhance semantic awareness, we provide the full score vector $\mathbf{s}_{i}$ of dataset-specific scores to the LLM in the solution generation stage, rather than averaging them into a single aggregate value. This enables the LLM to reason explicitly about behavioral differences across tasks.
3.5 Incorporating Domain Knowledge into Prompts
Algorithm evolution requires deep domain expertise, which general-purpose LLMs often lack. To compensate, domain knowledge is commonly incorporated into prompts to enhance LLM-based algorithm evolution [52, 35]. For the design of selection operators, we embed the following principles into the prompt to guide the LLM toward generating more effective solutions:
Diversity-aware: Purely objective-driven selection can cause premature convergence. To counter this, it is desirable to select models that perform well on different training instances to maintain population diversity.
Interpretability-aware: Interpretability, often measured by the number of nodes in a symbolic expression, is a critical criterion. Let $n_{i}$ denote the number of nodes in solution $p_{i}$. The selection operator should favor solutions with smaller $n_{i}$ to promote simpler, more interpretable models.
Stage-aware: Selection pressure should adapt over time. In early generations ($t\ll T_{\text{max}}$), it should favor exploration. In later stages ($t\approx T_{\text{max}}$), it should prioritize exploitation of high-fitness solutions to encourage convergence.
Complementarity-aware: To effectively recombine useful traits, crossover should favor solutions with complementary strengths. Given performance vectors $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ over $d$ instances, complementarity means selecting pairs with low correlation. Such combinations integrate distinct capabilities and can yield offspring that perform well across a broader range of instances.
Vectorization-aware: Vectorized operations are preferred, as they can be efficiently accelerated using NumPy or GPU computation. This improves the runtime efficiency of the evolved operator and reduces evaluation overhead during meta-evolution.
The prompt incorporating these domain knowledge principles is provided in Figure 15 of the supplementary material.
4 Experimental Settings
Meta-Evolution: The experiments are conducted using GPT-4.1 Mini. For the outer loop, the population size $N$ is set to 20, the number of solutions generated by mutation $M$ is set to 1, and the number of solutions generated by crossover is 19. The total number of generations is set to 20. A length limit of 30 lines is imposed to control operator complexity. For the inner loop, four datasets are selected to evaluate the quality of the generated selection operators. These datasets have OpenML IDs 505, 4544, 588, and 650, corresponding to the four highest-dimensional datasets in the contemporary symbolic regression benchmark (SRBench) [31]. High-dimensional datasets pose greater challenges for symbolic regression algorithms, and optimizing on these is likely to generalize well to both easy and hard problems. The number of generations in the inner loop is set to 30, and the population size is 100. More detailed experimental settings for the inner symbolic regression loop are provided in Appendix D of the supplementary material. All experiments are repeated three times with different random seeds [67] to report the median and confidence intervals.
Symbolic Regression: For symbolic regression, experiments are conducted on 116 out of the 120 datasets in SRBench, excluding the 4 datasets used during the meta-evolution phase to prevent potential data leakage. Each algorithm is evaluated using 10 independent runs per dataset to ensure statistically robust comparisons. The symbolic regression algorithm used is standard genetic programming with linear scaling, which is a widely adopted framework in symbolic regression research [61, 29]. The parameter settings are the same as those used in the inner symbolic regression loop, except that the number of generations is increased to 100.
5 Experimental Results
5.1 Meta-Evolution Results
In this section, we conduct a series of experiments to evaluate the performance of our proposed meta-evolution framework (LLM-Meta-SR). Specifically, we compare the full method against several ablated variants: without semantic evolution (W/O Semantics), without semantic evolution and bloat control (W/O SE+BC), without domain knowledge (W/O Knowledge), without domain knowledge and semantic evolution (W/O DK+SE), and without domain knowledge, semantic evolution, and bloat control (W/O DK+SE+BC). Since domain knowledge is not always available in practice, we include ablation studies under the setting without domain knowledge to evaluate how the proposed strategy performs in entirely new domains. In the W/O Semantics setting, random selection with identical objective prevention from ReEvo [69] is used as the baseline. In the W/O SE+BC setting, survival selection is replaced with direct population replacement and elitism, following the strategy used in ReEvo [69]. In the W/O Knowledge setting, semantic evolution and bloat control are retained, but the domain knowledge prompt is removed.
Objective Score: The objective score is measured by the average test $R^{2}$ score of each selection operator across four symbolic regression training datasets. For the W/O DK+SE+BC setting, only one run completed within 24 hours and is thus reported. The results in Figure 5, along with the numerical values presented in Table 1, show that the absence of domain knowledge results in the most significant performance drop. This highlights the importance of domain knowledge in guiding effective algorithm evolution, validating the principles introduced in Section 3.5. This performance gain is consistent with the findings of FunSearch [52] and EoH [35], which suggest that LLMs alone often fail to generate sufficiently good solutions without external guidance.
However, domain knowledge alone is insufficient to discover high-performing operators. As shown in Figure 5, semantic evolution plays a critical role in achieving higher objective scores, both with and without domain knowledge. To investigate the role of semantics, we visualize the semantic distribution of solutions over the first five generations in Figure 6. Each point represents a solution, colored by its average score across the four tasks. Positions are computed via t-SNE based on the four-dimensional score vector. The figure reveals that the number of solutions achieving top-3 performance on at least one dataset consistently exceeds three, indicating that top-performing individuals in terms of average score do not dominate across all tasks. Additionally, the final subplot in Figure 6 shows that some solutions, marked with stars, achieve top-3 performance on at least one dataset despite having modest average scores. This illustrates that relying solely on average performance may overlook individuals with strong task-specific capabilities. These findings highlight the importance of semantic-based algorithm evolution in capturing diverse and task-specific behaviors that are obscured by aggregate metrics.
Table 1: Historical best scores and corresponding code lengths for each algorithm.
LLM-Meta-SR W/O Semantics W/O SE+BC W/O Knowledge W/O DK+SE W/O DK+SE+BC
Score 0.86 0.84 0.84 0.79 0.75 0.75
Lines of Code 48 43 80 55 41 222
Figure 4: Validation $R^{2}$ of the best solution across generations for different LLM-driven search strategies.
Figure 5: Code length of the best solution across generations for different LLM-driven search strategies.
Figure 6: t-SNE visualization of evolved operator semantics. The shape of each point indicates whether it achieved top-3 performance on any dataset: stars denote top-3 performance on at least one dataset, while circles indicate otherwise.
Code Length: Figure 5 shows the evolution of code length for the best-performing solutions. Without bloat control, code length grows excessively throughout the optimization process when domain knowledge is absent. Incorporating domain knowledge helps shape the basic structure and mitigates excessive growth, but the code length still remains substantial. In contrast, when bloat control is applied, it directly curbs the complexity bias of LLMs, resulting in smaller generated programs that stabilize at a length of around 50, while simultaneously achieving faster improvements in objective scores.
5.2 Discovered Operators
In this section, we present experiments on datasets from contemporary symbolic regression benchmarks to evaluate the performance of the discovered operators. Since the LLM-designed operator incorporates several desirable characteristics of selection operators—including diversity-awareness and stage-awareness—it is referred to as Omni selection (Omni). The code for the Omni selection operator is provided in Appendix B of the supplementary material. We compare the performance of the LLM-generated selection operator against several expert-designed baselines, including automatic $\epsilon$-lexicase selection (AutoLex) [30], probabilistic lexicase selection (PLex) [13], lexicase-like selection via diverse aggregation (DALex) [45], $\epsilon$-lexicase selection with dynamic split (D-Split) [25], random sampling with tournament selection (RDS-Tour) [20], complementary phenotype selection (CPS) [14], tournament selection with sizes 3 and 7 [2], and Boltzmann selection with temperature scheduling (Boltzmann) [56]. All of these operators were manually designed by domain experts.
Figure 7: Test $R^{2}$ scores of different selection operators on symbolic regression benchmarks. Median values are annotated.
Figure 8: Pairwise statistical comparison of selection operators using the Wilcoxon signed-rank test with Benjamini–Hochberg correction.
Test Accuracy: The results in Figure 8 show that the LLM-generated selection operator outperforms the expert-designed baselines in terms of $R^{2}$ scores and achieves the best overall performance. A Wilcoxon signed-rank test with Benjamini–Hochberg correction is presented in Figure 8. These results demonstrate that the evolved Omni selection operator performs significantly better than existing expert-designed operators, many of which cannot be statistically distinguished from one another. This confirms that LLMs can effectively discover selection operators that surpass those created by domain experts.
Figure 9: Tree sizes of different selection operators on symbolic regression benchmarks.
Figure 10: Training times of different selection operators on symbolic regression benchmarks.
Model Size: The distribution of model sizes is presented in Figure 10. The results show that the evolved operator produces smaller models compared to top-performing selection operators like AutoLex and CPS. This is primarily because the evolved operator explicitly incorporates model size into the selection process, biasing the search toward regions where symbolic expressions are more compact. As shown in Figure 10 and Figure 8, model size and performance are not always in conflict. By carefully designing the selection operator, it is possible to evolve symbolic models that achieve both high accuracy and small model size, indicating high interpretability.
Training Time: As shown in Figure 10, the proposed Omni selection operator is efficient, even though it integrates multiple criteria into the selection process, which makes it slightly more time-consuming than some other operators in terms of overall symbolic regression time. However, it is important to consider not only the training phase but also the cost of evaluating and deploying candidate models. Since Omni selection tends to favor smaller models with fewer nodes, it reduces computational cost during evaluation and deployment, especially on resource-constrained devices. Thus, despite its multi-faceted design leading to a slight increase in selection time, Omni remains an efficient and practical selection operator.
6 Conclusions and Limitations
In this paper, we propose an LLM-driven framework for automatically designing selection operators in symbolic regression, and design several strategies to enhance its effectiveness. First, we highlight the challenges of bloat and limited semantic awareness in LLM-driven evolution. To mitigate bloat, we introduce a prompt-based length control and a multi-objective survival selection strategy. Moreover, we propose a semantic-based selection operator and a semantic feedback mechanism. Additionally, we define a set of design principles derived from domain knowledge to better guide LLM for operator generation. Ablation studies confirm that addressing these three factors significantly enhances the effectiveness of the evolved algorithms. Next, we evaluate the LLM-generated selection operators against expert-designed baselines across a wide range of symbolic regression datasets, demonstrating that LLMs can discover selection operators that outperform those crafted by domain experts.
The limitation of this paper is that, while the design of selection operators for symbolic regression is automated, the algorithmic components of the meta-evolution framework remain manually designed. A promising direction for future work is to develop a hyper-evolution framework that can automatically evolve the meta-evolution process itself. Another limitation is the reliance on handcrafted prompts incorporating domain knowledge. Future work may explore fully automated algorithm design by enabling LLMs to retrieve relevant online information and synthesize new algorithmic strategies autonomously.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims include a semantics-aware and bloat-controlled, LLM-based automated algorithm design framework for effectively designing a selection operator for symbolic regression. These claims accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The proposed method is described in detail in Section 3, and the experimental settings are provided in Section 4. Additionally, our code is publicly available, which further supports reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code is available, with sufficient instructions to faithfully reproduce the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Section 4 and Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: See Figure 5, Figure 5, and Figure 8.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Section D.5.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Automated algorithm design can help researchers develop more effective algorithms and reduce the time and effort required for manual design. The broader impacts of this work are discussed in Section 1.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: See Appendix I.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The paper describes the detailed usage of LLMs as a mechanism to automatically generate selection operators for symbolic regression.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Prompt for Algorithm Evolution
System Prompt: Figure 11 shows the system prompt used for algorithm evolution, which is added before the specific prompt in all three phases of the automated algorithm design process, including initialization, crossover, and mutation. The system prompt is designed to guide the LLM in generating efficient selection operators by preferring NumPy vectorized operations and avoiding explicit Python for-loops unless absolutely necessary. Python is chosen for its simplicity and strong support in LLMs, but standard for-loops are time-consuming. In contrast, NumPy vectorized operations are implemented in C++ and offer better computational speed. The system prompt also specifies the maximum number of lines of code that the LLM can generate. This constraint helps control bloat, as discussed in Section 3.3.
Initialization/Mutation Prompt: Figure 12 presents the prompts used for initialization and mutation. The main goal of these prompts is to encourage the LLM to generate a novel selection operator to explore the search space. The key difference between the prompts used for the two phases is that in the initialization phase, no baseline code is provided to the LLM. In the mutation phase, the elite solution is provided as the baseline code. The prompt format for wrapping the baseline code is shown in Figure 13, which aims to encourage the LLM to generate a novel selection operator based on the provided code.
Crossover Prompt: Figure 14 shows the crossover prompt used for algorithm evolution. The goal is specified as “selection operator,” and its plural form, “selection operators”. The crossover prompt takes the code of two existing selection operators as input and aims to generate a better operator by combining effective building blocks from both. To support semantic awareness and concise code generation, the prompt includes the score vector and the corresponding line of code. The properties describe the desired characteristics of the selection operator based on domain knowledge from Section 3.5. The specific prompt is shown in Figure 15. The template provides a function signature to ensure that the generated operator can be integrated into the automatic evaluation pipeline as shown in LABEL:lst:template. When domain knowledge is not provided, the properties are left empty and a simplified template without knowledge guidance for the selection operator, as shown in LABEL:lst:template_without_domain_knowledge, is used.
Fallback Selection Operator: In rare cases where the LLM fails to generate valid Python code, a simple tournament selection operator, as shown in LABEL:lst:_tournament, is used to fill the population.
When writing code, prefer NumPy vectorized operations and avoid explicit Python for-loops unless absolutely necessary. Please implement code within {max_code_lines} lines.
Figure 11: System prompt for evolving selection operators.
Your task is to develop an innovative and novel {goal} for symbolic regression using genetic programming in Python. {Baseline} {Properties} Ensure that your newly designed function adheres to the following signature: {Template} You do not need to provide a usage example. Embrace creativity, novelty, and bold experimentation to push the boundaries of the state of the art in {goals} for genetic programming.
Figure 12: Initialization and mutation prompt for designing an LLM-based selection operator. For initialization, the baseline is empty.
Inspirational Example: {Code} Use this as inspiration to create a distinctly original and inventive {goal}.
Figure 13: Prompt embedded in the baseline.
You are tasked with designing a novel {goal} for symbolic regression using genetic programming. Your goal is to synthesize a new operator that combines the strengths and mitigates the weaknesses of the two operators shown below: — Operator A — Score (Higher is Better): {RoundToDecimalPlaces(Score_A, 3)}, Lines of Code: {EffectiveLineCount(OperatorA)} Code: {OperatorA_Code} — Operator B — Score (Higher is Better): {RoundToDecimalPlaces(Score_B, 3)}, Lines of Code: {EffectiveLineCount(OperatorB)} Code: {OperatorB_Code} {Properties} Ensure that your newly designed function adheres to the following signature: {Template} You do not need to provide a usage example.
Figure 14: Prompt used for crossover to design a novel selection operator with an LLM.
1. Diverse & Specialized Selection
     – Choose a varied set of high-performing individuals.
– Encourage specialization to maintain a diverse population. 2. Crossover-Aware Pairing
     – Promote complementarity between parents.
3. Stage-Specific Pressure
     – Vary selection pressure based on the current stage of evolution.
4. Interpretability
     – Prefer individuals with fewer nodes and lower tree height to improve model interpretability.
5. Efficiency & Scalability
     – Include clear stopping conditions to avoid infinite loops.
6. Code Simplicity
     – Favor clear, concise logic with minimal complexity.
Figure 15: Desired properties of good selection operators based on domain knowledge.
1def selection(population, k=100, status={}):
2 # Useful information about individuals:
3 # squared_error_vector = individual.case_values
4 # predicted_values = individual.predicted_values
5 # residual = individual.y-individual.predicted_values
6 # number_of_nodes = len(individual)
7 # height = individual.height
8
9 # Useful information about evolution:
10 # status["evolutionary_stage"]: [0,1], where 0 is the first generation and 1 is the final generation
11
12 parent_a = Select k//2 individuals based on performance over subsets of instances:
13 - evaluate individuals on k//2 different random or structured subsets of the data
14 - reward those that specialize or perform well on those subsets
15 - may also consider overall fitness (e.g., full-dataset MSE)
16 - optionally include structural complexity
17 - an individual can be selected multiple times
18
19 parent_b = For each parent_a, select a complementary individual:
20 - low residual correlation with parent_a
21 - may also consider complexity
22
23 # Interleave parent_a and parent_b to form crossover pairs
24 selected_individuals = [individual for pair in zip(parent_a, parent_b) for individual in pair]
25 return selected_individuals
Code 1: Template for selection operator evolution.
1def selection(population, k=100, status={}):
2 # Useful information about individuals:
3 # squared_error_vector = individual.case_values
4 # predicted_values = individual.predicted_values
5 # residual = individual.y-individual.predicted_values
6 # number_of_nodes = len(individual)
7 # height = individual.height
8
9 # Useful information about evolution:
10 # status["evolutionary_stage"]: [0,1], where 0 is the first generation and 1 is the final generation
11
12 # Implement selection logic here
13 return selected_individuals
Code 2: Template for selection operator evolution without domain knowledge.
1import numpy as np
2import random
3
4def selection(individuals, k=100, tour_size=3):
5 # Tournament selection
6 # Compute mean error for each individual
7 mean_errors = [np.mean(ind.case_values) for ind in individuals]
8 selected_individuals = []
9
10 # Select ‘k‘ individuals
11 for _ in range(k):
12 # Randomly select competitors
13 competitors = random.sample(range(len(individuals)), tour_size)
14 # Select the one with the lowest error
15 best_idx = min(competitors, key=lambda idx: mean_errors[idx])
16 selected_individuals.append(individuals[best_idx])
17
18 return selected_individuals
Code 3: Default code used when no code can be extracted from the LLM response.
Appendix B Analysis of an Evolved Selection Operator
B.1 Code Evolved by LLM-Meta-SR
B.1.1 Static Analysis
LABEL:lst:omni-selection shows the best evolved selection operator among all runs. The code has been simplified by removing logically redundant parts. In selecting the first parent (parent_a), the operator evaluates two key criteria: specificity—quantified as the lowest error achieved on any subset of the dataset, with lower values indicating superior performance in specialized regions—and a complexity measure that jointly considers the number of nodes in the symbolic expression tree and the height of the tree, thereby reflecting the interpretability of the solution. Candidate solutions are ranked such that subset error is prioritized, with model complexity serving as a secondary criterion when the subset error alone does not sufficiently differentiate candidates. This approach encourages the identification of solutions that balance high accuracy with interpretability.
For the selection of the second parent (parent_b), the operator computes the absolute cosine similarity between the residuals of each candidate in the population and the residual of the previously selected first parent. The second parent is then chosen based on a combination of semantic complementarity, as measured by cosine similarity, and model complexity. The trade-off between complementarity and complexity is controlled by a linear function related to the evolutionary stage. In the early stages, the pressure toward parsimony is intentionally relaxed to promote broader exploration of the search space. In contrast, in later stages, the pressure increases to favor models that are both accurate and concise, thus facilitating interpretability.
Overall, the selection operator satisfies all the desired properties, including interpretability, diversity, complementarity, stage-awareness, and vectorization, as defined in Section 3.5.
1def omni_selection(population, k=100, status={}):
2 stage = np.clip(status.get("evolutionary_stage", 0), 0, 1)
3 n = len(population)
4 half_k = k // 2
5 y = population[0].y
6 n_cases = y.size
7
8 ssize = max(7, n_cases // max(1, 2 * half_k))
9 half_struct = half_k // 2
10 structured = [
11 np.arange(i * ssize, min((i + 1) * ssize, n_cases)) for i in range(half_struct)
12 ]
13 random_ = [
14 np.random.choice(n_cases, ssize, replace=False)
15 for _ in range(half_k - half_struct)
16 ]
17 subsets = structured + random_
18
19 residuals = np.vstack([ind.y - ind.predicted_values for ind in population])
20 complexity = np.array([len(ind) + ind.height for ind in population], float)
21 complexity /= max(1, complexity.max())
22 subset_mse = np.array(
23 [
24 [((residuals[i, s]) ** 2).mean() if s.size else np.inf for s in subsets]
25 for i in range(n)
26 ]
27 )
28 comp_factor = 0.25 + 0.25 * stage
29
30 parent_a = [
31 population[np.lexsort((complexity, subset_mse[:, i]))[0]]
32 for i in range(len(subsets))
33 ]
34
35 idx_map = {ind: i for i, ind in enumerate(population)}
36 norms = np.linalg.norm(residuals, axis=1) + 1e-12
37
38 parent_b = []
39 for a in parent_a:
40 i_a = idx_map[a]
41 res_a = residuals[i_a]
42 cors = (residuals @ res_a) / (norms * norms[i_a])
43 cors[i_a] = 1
44 comp_score = np.abs(cors) + comp_factor * complexity
45 b_idx = np.argmin(comp_score)
46 parent_b.append(population[b_idx])
47
48 return [ind for pair in zip(parent_a, parent_b) for ind in pair][:k]
Code 4: Omni Selection
Figure 16: Test $R^{2}$ scores across generations.
Figure 17: Population cosine semantic diversity across generations.
Figure 18: Tree sizes across generations.
B.1.2 Benchmark Analysis
To further analyze the behavior of the Omni selection operator, we plot the test $R^{2}$ scores on five representative datasets in Figure 16. The diversity of the population is shown in Figure 17, and the tree size is shown in Figure 18.
Test $R^{2}$ Scores: The test $R^{2}$ scores demonstrate that the Omni selection operator achieves an advantage at the beginning of the search, indicating that it is an effective selection operator for anytime performance [68].
Population Diversity: To understand why the Omni selection operator achieves this performance, we plot the population diversity trajectory in Figure 17. Diversity is measured by the cosine distance between individuals, which is preferred over Euclidean distance because the latter can be trivially large if some solutions have very large errors. The results show that population diversity is well maintained across generations and is better than the baseline. In other words, the Omni selection operator maintains a more diverse population, with individuals making errors on different subsets of instances. This helps explain the superior performance of the Omni selection operator compared to the baseline.
Tree Size: Finally, we plot the tree size trajectory in Figure 18. The results show that Omni selection not only achieves better performance but also suffers less from rapid growth in tree size compared to other selection operators throughout the entire evolution process. This suggests that the Omni selection operator explores regions of the search space containing parsimonious solutions more exhaustively than other operators, which is an ideal behavior for finding effective and parsimonious solutions in symbolic regression.
B.2 Code Evolved by LLM-Meta-SR without Domain Knowledge
In this section, we analyze the code evolved by LLM-Meta-SR without domain knowledge to understand how an LLM can generate algorithms based solely on its internal knowledge. The evolved code is shown in LABEL:lst:omni-selection-free. We refer to the resulting algorithm as Omni-Zero, as it evolves the model from scratch, similar to AutoML-Zero [real2020automl].
B.2.1 Static Analysis
Based on the code in LABEL:lst:omni-selection-free, the LLM can evolve selection operators with desirable properties even without domain-specific guidance. These properties include diversity-awareness, stage-awareness, and interpretability-awareness. For diversity-awareness, the selection operator uses cosine similarity to measure pairwise distances between individuals. Formally, it defines the novelty score as:
$\mathrm{nov}_{i}=\frac{1-\frac{1}{n}\sum_{k=1}^{n}\left\langle\frac{\mathbf{x}_{i}-\bar{x}_{i}}{\|\mathbf{x}_{i}-\bar{x}_{i}\|_{2}+\epsilon},\frac{\mathbf{x}_{k}-\bar{x}_{k}}{\|\mathbf{x}_{k}-\bar{x}_{k}\|_{2}+\epsilon}\right\rangle}{\max\limits_{j}\left[1-\frac{1}{n}\sum_{k=1}^{n}\left\langle\frac{\mathbf{x}_{j}-\bar{x}_{j}}{\|\mathbf{x}_{j}-\bar{x}_{j}\|_{2}+\epsilon},\frac{\mathbf{x}_{k}-\bar{x}_{k}}{\|\mathbf{x}_{k}-\bar{x}_{k}\|_{2}+\epsilon}\right\rangle\right]}$ (3)
This cosine-distance-based novelty score is better suited for symbolic regression than Euclidean-distance-based novelty scores, as it is more robust to outliers and cannot be easily manipulated by increasing the distance with some outlier, as discussed in Section B.1.2.
For stage-awareness, the selection operator applies a non-linear weighting function to control the trade-off between error and novelty. The function is defined as $0.35+0.3\left(1-\frac{t}{T_{\max}}\right)^{0.8}$, where $t$ is the current generation and $T_{\max}$ is the maximum number of generations, as described in LABEL:lst:omni-selection-free. This function gradually decreases over time, promoting exploration in early generations and exploitation in later ones. This dynamic trade-off satisfies the stage-awareness criterion defined in Section 3.5.
For interpretability-awareness, the selection operator uses both the number of nodes and the height of the tree to measure solution complexity. A linear weighting function is used to balance interpretability and accuracy during selection.
A limitation of this operator is that it does not consider diversity from the perspective of specificity, the selection of each individual is largely dominated by the mean squared error of each individual, nor does it account for the complementarity between two parents.
1def omni_selection(population, k=100, status={}):
2 import numpy as np
3 evo_stage = status.get("evolutionary_stage", 0.)
4 n = len(population)
5 if n == 0:
6 return []
7
8 # Extract metrics
9 errs = np.array([np.mean(ind.case_values) for ind in population], dtype=float)
10 sizes = np.array([len(ind) for ind in population], dtype=float)
11 heights = np.array([ind.height for ind in population], dtype=float)
12 residuals = np.array([ind.y - ind.predicted_values for ind in population])
13 preds = np.array([ind.predicted_values for ind in population])
14
15 def safe_norm(arr):
16 m = max(arr.max(), 1e-10)
17 return arr / m
18
19 norm_size, norm_height = safe_norm(sizes), safe_norm(heights)
20 res_vars = np.var(residuals, axis=1)
21 res_norms = np.linalg.norm(residuals, axis=1)
22 norm_var, norm_resnorm = safe_norm(res_vars), safe_norm(res_norms)
23 norm_err = safe_norm(errs)
24
25 # Complexity metric combines more residual stats while balancing structural features
26 complexity = (norm_size + norm_height + norm_var + norm_resnorm) / 4
27 alpha, beta = 1 - evo_stage, evo_stage
28 base_score = beta * (1 - norm_err) + alpha * (1 - complexity)
29 base_score -= base_score.min()
30 if base_score.sum() == 0:
31 base_score[:] = 1
32 base_probs = base_score / base_score.sum()
33
34 # Novelty score based on residuals and predicted values diversity
35 def novelty_score(mat):
36 centered = mat - mat.mean(axis=1, keepdims=True)
37 norms = np.linalg.norm(centered, axis=1, keepdims=True) + 1e-10
38 normed = centered / norms
39 sim = normed @ normed.T
40 nov = 1 - sim.mean(axis=1)
41 max_n = nov.max()
42 return nov / (max_n if max_n > 0 else 1)
43
44 novelty_res = novelty_score(residuals)
45 novelty_pred = novelty_score(preds)
46 novelty = 0.5 * (novelty_res + novelty_pred)
47
48 # Dynamic novelty weighting: stronger early novelty, moderate late novelty
49 novelty_weight = 0.35 + 0.3 * (1 - evo_stage)**0.8
50
51 mixed_probs = (1 - novelty_weight) * base_probs + novelty_weight * novelty
52 mixed_probs -= mixed_probs.min()
53 if mixed_probs.sum() == 0:
54 mixed_probs[:] = 1
55 mixed_probs /= mixed_probs.sum()
56
57 # Adaptive tournament size encourages pressure on error, preserves diversity
58 tour_size = min(3 + int(evo_stage * 2), n)
59 selected = []
60 while len(selected) < k:
61 chosen = np.random.choice(n, size=tour_size, replace=False, p=mixed_probs)
62 # Tournament winner: lowest error considering base scores and novelty tie-break
63 chosen_errs = errs[chosen]
64 chosen_scores = base_probs[chosen]
65 min_err_idx = chosen_errs.argmin()
66 # Tie-break with highest combined score
67 best_candidates = np.flatnonzero(chosen_errs == chosen_errs[min_err_idx])
68 if len(best_candidates) > 1:
69 best_scores = chosen_scores[best_candidates] + novelty[chosen][best_candidates]
70 winner_idx = best_candidates[np.argmax(best_scores)]
71 else:
72 winner_idx = min_err_idx
73 selected.append(population[chosen[winner_idx]])
74 return selected
Code 5: Omni Selection Zero
B.2.2 Benchmark Analysis
The Omni-Zero operator is evaluated on the SRBench datasets [la2021contemporary], and the results are shown in Figure 20. As shown by the results, the proposed method achieves medium performance among the benchmarked algorithms. It performs significantly better than tournament selection, as indicated by the p-values in Figure 20. However, compared to selection operators that consider specificity, such as the lexicase selection operator, the evolved selection operator is inferior.
Figure 19: Test $R^{2}$ scores of different selection operators on symbolic regression benchmarks, including Omni-Zero.
Figure 20: Pairwise statistical comparison of selection operators using the Wilcoxon signed-rank test with Benjamini–Hochberg correction.
Figure 21: Tree sizes of different selection operators on symbolic regression benchmarks, including Omni-Zero.
Figure 22: Training times of different selection operators on symbolic regression benchmarks, including Omni-Zero.
B.3 A Repaired Version of Omni Selection for Small Datasets
B.3.1 Problem Analysis
Upon closely analyzing the code generated by the LLM in LABEL:lst:omni-selection, we identified a logical bug that can lead to poor performance of the selection operator on small datasets. When the dataset is small, the structured division in line 10 of LABEL:lst:omni-selection does not yield enough samples to form sufficient subsets, which can result in some subsets being empty. These empty subsets cause the subsequent selection process to rely solely on complexity when selecting individuals, without considering their predictive performance. This issue arises because the datasets used during meta-evolution generally contain a relatively large number of training instances, as shown in Table 2. Consequently, the bug has limited impact during training but becomes problematic when applied to smaller datasets.
Table 2: Summary of datasets used in meta-evolution.
Dataset n_observations n_features
OPENML_505 240 124
OPENML_4544 1059 117
OPENML_588 1000 100
OPENML_650 500 50
B.3.2 Solution
To address this issue, we propose a repaired version of the Omni selection operator, as shown in LABEL:lst:omni-selection-repaired. In this version, random subsets are used to replace any empty subsets generated during structured division. All other parts of the code remain unchanged.
B.3.3 Benchmark Analysis
We compare the performance of the repaired and original versions on datasets with no more than 100 training instances. The results, presented in Figure 24, show that the repaired version outperforms the original, indicating that the logical bug in the original version is indeed problematic. Regarding model size, as shown in Figure 24, the repaired version results in an increased tree size, which is reasonable because it corrects the previous behavior of selecting individuals solely based on complexity by also considering the error on subsets. Nonetheless, the increase in tree size is not substantial and still leads to a competitive model size compared to other selection operators.
Figure 23: Test $R^{2}$ scores of different selection operators on symbolic regression benchmarks, including the repaired omni selection.
Figure 24: Tree sizes of different selection operators on symbolic regression benchmarks, including the repaired omni selection.
This failure highlights that, when using LLMs for meta-evolution, it is important that the LLM is exposed to a wide range of instances. Otherwise, the generated code may contain subtle logical bugs that are not easily detected.
1def omni_selection(population, k=100, status={}):
2 stage = np.clip(status.get("evolutionary_stage", 0), 0, 1)
3 n = len(population)
4 half_k = k // 2
5 y = population[0].y
6 n_cases = y.size
7
8 ssize = max(7, n_cases // max(1, 2 * half_k))
9 half_struct = half_k // 2
10 structured = [
11 np.arange(i * ssize, min((i + 1) * ssize, n_cases)) for i in range(half_struct)
12 ]
13 # Filter out empty subsets
14 structured = [s for s in structured if s.size > 0]
15 random_ = [
16 np.random.choice(n_cases, ssize, replace=False)
17 for _ in range(half_k - len(structured))
18 ]
19 subsets = structured + random_
20
21 residuals = np.vstack([ind.y - ind.predicted_values for ind in population])
22 complexity = np.array([len(ind) + ind.height for ind in population], float)
23 complexity /= max(1, complexity.max())
24 subset_mse = np.array(
25 [
26 [((residuals[i, s]) ** 2).mean() if s.size else np.inf for s in subsets]
27 for i in range(n)
28 ]
29 )
30 comp_factor = 0.25 + 0.25 * stage
31
32 parent_a = [
33 population[np.lexsort((complexity, subset_mse[:, i]))[0]]
34 for i in range(len(subsets))
35 ]
36
37 idx_map = {ind: i for i, ind in enumerate(population)}
38 norms = np.linalg.norm(residuals, axis=1) + 1e-12
39
40 parent_b = []
41 for a in parent_a:
42 i_a = idx_map[a]
43 res_a = residuals[i_a]
44 cors = (residuals @ res_a) / (norms * norms[i_a])
45 cors[i_a] = 1
46 comp_score = np.abs(cors) + comp_factor * complexity
47 b_idx = np.argmin(comp_score)
48 parent_b.append(population[b_idx])
49
50 return [ind for pair in zip(parent_a, parent_b) for ind in pair][:k]
Code 6: The Repaired Version of Omni Selection (Omni-R)
Appendix C Test Cases for Synthetic Evaluation
For the synthetic evaluation in Section 3.2, two synthetic test cases are designed to reveal implementation flaws and inefficiencies early in the evaluation process, prior to applying each evolved selection operator to real symbolic regression tasks. The first case consists of 100 solutions whose predicted values and training errors are independently sampled from 100 random integers in the range $[1,10]$. The second case also consists of 100 solutions, but with both predicted values and training errors set to a constant integer, repeated across all individuals, based on a value sampled from $[1,10]$. The first case represents a scenario with diverse inputs and varying fitness values, while the second serves as an edge case where all individuals have identical fitness.
Appendix D Evolutionary Symbolic Regression
D.1 Solution Representation
The symbolic regression model is represented as a symbolic expression tree with linear scaling [nadizar2024geometric]. For a given expression $\Phi$, the prediction is computed as $\hat{Y}=\alpha\cdot\Phi(X)+\beta$, where the coefficients $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{R}$ are fitted using ridge regression by minimizing the regularized squared error between $\hat{Y}$ and the targets $Y$.
D.2 Algorithm Workflow
The symbolic regression algorithm is implemented using the genetic programming framework [banzhaf1998genetic]. Let $\mathcal{F}^{(t)}=\{\Phi_{1},\Phi_{2},\dots,\Phi_{\lambda}\}$ denote the population of $\lambda$ symbolic expressions at generation $t$. The evolutionary process proceeds as follows:
  • Population Initialization: The initial population $\mathcal{F}^{(0)}$ is generated using the ramped half-and-half method [banzhaf1998genetic].
[…]
  • Elitism: The best-performing expression $\Phi^{*}=\arg\min_{\Phi_{i}\in\mathcal{F}^{(t)}}\mathcal{E}_{i}$ is preserved in an external archive and used for solution selection to maintain historical performance.
  • Solution Selection: The LLM-evolved selection operator is invoked to select a set of promising parent expressions from the current population and the elite archive.
  • Solution Generation: A new population $\mathcal{F}^{(t+1)}$ is generated from the selected parents using standard genetic programming operators:
    • Subtree Crossover: Given two parent expressions $\Phi_{a}$ and $\Phi_{b}$, offspring are generated by exchanging randomly selected subtrees.
    • Subtree Mutation: A randomly selected subtree of an expression $\Phi_{i}$ is replaced with a newly generated subtree.
    During solution generation, all solutions generated in the evolutionary process are stored in a hash set. If an offspring is identical to any historical solution, it is discarded and a new solution is generated. The hash set is used because it maintains $O(1)$ query complexity. This mechanism prevents symbolic regression from wasting resources by evaluating the same solution multiple times.
D.3 Parameter Settings
The parameters for symbolic regression follow conventional settings. The population size and number of generations are set to match those used in D-Split [imai2024minimum]. To prevent division-by-zero errors, we use the analytical quotient ($\mathrm{AQ}(x,y)=\frac{x}{\sqrt{1+y^{2}}}$[ni2012use] in place of the standard division operator.
Table 3: Parameter settings
Parameter Value
Population size 100
Maximum generations 100
Maximum tree depth 10
Initialization method Ramped half-and-half (depth 0-6)
Function set +,,×$+,\,-,\,\times,\,\mathrm{AQ},\,\sqrt{\cdot},\,\log(1+|\cdot|),\,|\cdot|,\,(\cdot)^{2},$ $\sin_{\pi}(\cdot),\,\cos_{\pi}(\cdot),\,\mathrm{Max},\,\mathrm{Min},\,\mathrm{Neg}$
Crossover rate 0.9
Mutation rate 0.1
D.4 Evaluation Protocol in Meta-Evolution
In the evaluation of selection operators generated by the LLM, all other algorithmic components—including solution initialization, evaluation, generation, and elitism—follow standard practices commonly used in evolutionary symbolic regression. For each dataset $\mathcal{D}_{j}$, the data is split into training and validation subsets with an 80:20 ratio [la2021contemporary]. The regression model is optimized to maximize the training $R^{2}$ score, while the validation $R^{2}$ score is used as the final score of the selection operator. This encourages the selection of operators with implicit regularization capabilities.
D.5 Evaluation Protocol for Discovered Operators
For the experiment evaluating the evolved selection operator on SRBench, the dataset is randomly split into a training set and a test set with an 80:20 ratio [la2021contemporary]. Subsampling is applied to limit the maximum number of training instances to 10,000, consistent with the SRBench protocol [la2021contemporary]. Symbolic regression is run on the training set and evaluated on the test set. All features are standardized, as this has been shown to be beneficial for symbolic regression [owen2022standardization]. Standardization parameters are learned from the training data and applied to both the training and test sets. All symbolic regression experiments are conducted on an AMD Milan 7713 CPU.
Appendix E Baseline Selection Operators
The hyperparameters for the baseline selection operators follow the default settings specified in their respective original papers. For RDS-Tour, the sampling ratio is set to 10% and the tournament size to 7 [geiger2025tournament]. For PLex selection, the temperature parameter is set to 1.0 [ding2023probabilistic]. For DALex, a particularity pressure of 3 is used [ni2024dalex]. For CPS, the first parent is selected via tournament, with the tournament size set to 7 [xu2022genetic].
For Boltzmann sampling with temperature scheduling [shojaee2025llmsr, romera2024mathematical], a temperature decay rule is applied to balance exploration and exploitation over time, similar to the approaches used in LLM-SR [shojaee2025llmsr] and FunSearch [romera2024mathematical]. The sampling probability for individual $i$ is defined as:
$P_{i}=\frac{\exp\left(\frac{s_{i}}{\tau_{c}}\right)}{\sum_{i^{\prime}}\exp\left(\frac{s_{i^{\prime}}}{\tau_{c}}\right)},\quad\tau_{c}=\tau_{0}\left(1-\frac{t\bmod T_{\max}}{T_{\max}}\right),$ (4)
where $s_{i}$ denotes the score of individual $i$, $\tau_{0}=0.1$ is the initial temperature [shojaee2025llmsr], $t$ is the current generation, and $T_{\max}$ is the total number of generations. This scheduling rule linearly decays the temperature over generations and gradually shifts the selection pressure from exploration (high temperature) to exploitation (low temperature) as evolution progresses in evolutionary symbolic regression.
Appendix F Further Investigation on Crossover and Mutation Ratio
In the main paper, the number of individuals in the population generated by crossover is 19, and the number generated by mutation is 1. In this section, we investigate the proposed method under different settings of crossover and mutation ratios to examine how these ratios affect its performance. Specifically, in this section, the number of individuals generated by crossover is 15 and by mutation is 5. The results related to validation score and code length are shown in Figure 26 and Figure 26, respectively.
First, the experimental results in Figure 26 and Figure 26 show that the conclusions from the ablation studies in Section 5.2 generalize to different settings of crossover and mutation ratios. This includes the effectiveness of semantics-based evolution, bloat control, and domain-knowledge-based operators.
Second, the results in Figure 26 as well as the numerical values in Table 4 indicate that the optimal crossover and mutation ratios vary depending on the prompts. When domain knowledge is available, a smaller mutation rate, as used in Section 5.2, achieves slightly better performance than the larger mutation rate used in this section. In contrast, when domain knowledge is not available, a larger mutation rate appears to be more beneficial. One possible reason is that when domain knowledge is available, the search space is more confined and a large mutation rate does not provide additional benefit. Conversely, when domain knowledge is not available, the search space is larger and a higher mutation rate is needed to effectively explore the space.
Table 4: Historical best scores and corresponding code lengths for each algorithm in the case of large mutation rate.
LLM-Meta-SR W/O Semantics W/O SE+BC W/O Knowledge
Score 0.84 0.82 0.82 0.83
Lines of Code 64 48 101 64
Figure 25: Validation $R^{2}$ of the best solution across generations for different LLM-driven search strategies in the case of large mutation rate.
Figure 26: Code length of the best solution across generations for different LLM-driven search strategies in the case of large mutation rate.
Appendix G Analysis of Token Count of Generated Code
The token count directly impacts the cost of algorithm evolution, as language service providers typically charge based on the number of tokens. In this section, we analyze how bloat influences the token count of generated code. Token counting is performed using the cl100k_base tokenizer, which is employed by OpenAI models. The experimental results are shown in Figure 27. It is evident that when no bloat control strategy is applied, the token count of the generated code is significantly higher. In contrast, the application of a bloat control strategy substantially reduces the token count in both the LLM-Meta-SR and LLM-Meta-SR without domain knowledge settings. This indicates that the bloat control strategy is not only helpful for improving the interpretability of generated code, but also beneficial for reducing the cost of algorithm evolution.
Figure 27: Average token count of generated code across generations for different LLM-driven search strategies.
Appendix H Analysis on State-of-the-Art Symbolic Regression Algorithms
H.1 Experimental Settings
In this section, we apply the evolved operator to a state-of-the-art symbolic regression algorithm, namely retrieval-augmentation-generation-based symbolic regression (RAG-SR) [zhang2025ragsr], to evaluate the effectiveness of the proposed operator in the context of modern symbolic regression. The only modification made is the replacement of the automatic epsilon lexicase selection [la2019probabilistic] with the repaired Omni selection operator described in LABEL:lst:omni-selection-repaired. All other components remain unchanged. The resulting algorithm is referred to as RAG-SR-Omni. For the experimental datasets, following the setup in Section 4, we use 116 out of 120 regression problems, excluding the four used during meta-evolution.
H.2 Experimental Results
The results are presented in Figure 28, with the top-5 performing algorithms shown separately in Figure 29 for better clarity. These figures demonstrate that RAG-SR-Omni outperforms RAG-SR in terms of median $R^{2}$ scores, model sizes, and training time. Furthermore, as shown by the Pareto front of the ranks of test $R^{2}$ scores and model sizes across different algorithms in Figure 30, RAG-SR-Omni establishes a new Pareto front on SRBench. Collectively, these results support the effectiveness of the LLM-evolved selection operator.
Recall that RAG-SR uses automatic-lexicase selection, which only considers diversity-awareness. These results suggest that even for a state-of-the-art symbolic regression algorithm, it is still desirable to consider multiple aspects, such as diversity, stage-awareness, complementarity, and interpretability, during selection to achieve stronger performance. Nonetheless, manually designing such an operator is challenging, which highlights the value of leveraging LLMs in this context.
Figure 28: Median $R^{2}$ scores, model sizes, and training time of 26 algorithms on the symbolic regression benchmark [la2021contemporary].
Figure 29: Median $R^{2}$ scores, model sizes, and training time of top-5 algorithms on the symbolic regression benchmark.
Figure 30: Pareto front of the rank of test $R^{2}$ scores and model sizes for different algorithms.
Appendix I Licenses for used assets
The licenses for the used assets are shown in Table 5.
Table 5: Used assets and their licenses.
Type Asset License Usage
Code DEAP [fortin2012deap] LGPL-3.0 License Evaluation
Code CodeBLEU [ren2020codebleu] MIT License Evaluation
Dataset SRBench [la2021contemporary] GPL-3.0 License Testing
Appendix J Related Work
J.1 Recent Advancements in Selection Operators
In recent years, lexicase selection has shown strong empirical performance and is widely adopted in modern symbolic regression systems [cava2018learning, zhang2025ragsr]. Inspired by lexicase selection, several variants have been developed, including probabilistic lexicase selection [ding2023probabilistic], lexicase-like selection via diverse aggregation (DALex) [ni2024dalex], $\epsilon$-lexicase selection with dynamic split (D-Split) [imai2024minimum], down-sampled lexicase selection [boldi2024informed], and random down-sampled tournament selection [geigerperformance]. These methods share the core idea of emphasizing performance on subsets of training instances to encourage specialization, and they demonstrate advantages in various scenarios [geigerperformance]. However, these operators are manually designed and do not fully satisfy all the desired properties outlined in Section 3.5, highlighting the need for further exploration of selection operators.
J.2 Automated Evolutionary Algorithm Design
Before the LLM era, genetic programming had been widely used for automated evolutionary algorithm design, including the design of fitness functions [fong2024metasr], parameter adaptation techniques [stanovov2022automatic], and update strategies [lones2021evolving]. More recently, reinforcement learning has also been applied to designing update rules in differential evolution [chen2024symbol]. However, these approaches are typically limited to generating only small code segments, as exploring large program spaces remains challenging. The emergence of LLMs has made automated algorithm design more feasible. For example, LLMs have been used to design a differential evolution algorithm competitive with state-of-the-art continuous optimization methods [van2024llamea]. Similarly, LLMs have been employed to improve the self-organizing migrating algorithm [pluhacek2024using]. This growing interest in LLMs motivates the development of efficient LLM-based approaches for automated algorithm design, where semantic awareness and code bloat are two key issues that remain underexplored and warrant further investigation.
J.3 Large Language Models for Optimization
LLMs have been employed as optimization tools across various domains [meyerson2024language, song2024position], including neural architecture search [chen2023evoprompting], graph neural network design [mo2025autosgnn], reward function design [maeureka], the traveling salesman problem [liu2024large], prompt optimization [yanglarge], automated machine learning [chan2024mle], solver code generation [JiangShu2025llmopt], symbolic regression [shojaee2025llmsr], and feature engineering [gong2025evolutionary]. Evolutionary search has gained considerable attention in the LLM era due to its effectiveness over random sampling [zhang2024understanding]. Various efforts have aimed to enhance LLM-based evolutionary search, including group relative policy optimization [huang2025calm], direct preference optimization [vsurinaalgorithm], core abstraction prompting [wu2025efficient], quality-uncertainty optimization [chen2024qube], and bi-dimensional co-evolution [zhao2025understanding].
Exploring the potential of LLMs in automating the design of improved evolutionary search strategies is a worthwhile direction. However, evaluating LLM-based evolutionary search in automated algorithm design is computationally expensive. In comparison, genetic programming is a relatively inexpensive and historically dominant method for code generation prior to the LLM era [peabody2015gef, chen2023symbolic, maudet2025search]. Therefore, leveraging LLMs for the automated design of genetic programming algorithms presents a relevant and cost-effective avenue for investigation.